If a sample of oxygen gas, originally at 5.00 atm in 1.00 L, is heated in that container from 25.0°C to 175°C, what is the final pressure? A: 7.52 atm B: 0.714 atm C: 0.133 atm D: 35.0 atm
Answer #1

P1 = 5.0 atm

A: 7.52 atm

V1 = 1.0 L

T1 = 25 oC = 298.15 K

P2 = to be calculated

T2 = 175 oC = 448.15 K

Volume (V2) = constant = 1 L

So,

P1 / T1 = P2 / T2

final Pressure (P2) = (P1 / T1) x T2 = (5.0 atm / 298.15 K) x (448.15 K) = 7.52 atm